OpenStudy (anonymous):

Company A rents copiers for a monthly charge of $200 plus 10 cents per copy. Company B rents copiers for a monthly charge of $400 plus 5 cents per copy. What is the number of copies above which Company A's charges are the higher of the two?

OpenStudy (anonymous):

\[P_A=0.1c+200\] \[P_B=0.05c+400\] We want to know for what 'c' \[P_A \geq P_B\] but we can just solve \[P_A = P_B\] so \[0.1c+200=0.05c+400\] and solve for 'c' (this is up to you).